artlover03:

A car is traveling at a rate of 30 miles per hour. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 2 hours? In your computations, assume that 1 mile is equal to 1.6 kilometers. Do not round your answers. rate= _ km/h distance traveled in 2 hours: _km

justus:

Did someone say cars? jk lol 1 mile is 1.6 km The rate is 30 km The rate (30 x 1) / 1.6 =

artlover03:

err... 18.75 or 75/4 ????

justus:

yes 18.75

justus:

now 18.75 x 2 =

jhonyy9:

@justus not 30 km there are 30 miles /h

justus:

oh yes my mistake sorry

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus now 18.75 x 2 = \(\color{#0cbb34}{\text{End of Quote}}\) you will get the distance after calculating this

artlover03:

37.5...?

justus:

yes

artlover03:

Awesome thank you!

justus:

not a problem

justus:

hmm do you have to round?

artlover03:

No sir :)

justus:

got it

justus:

good luck ! (:

jhonyy9:

A car is traveling at a rate of 30 miles per hour. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 2 hours? In your computations, assume that 1 mile is equal to 1.6 kilometers. Do not round your answers. 30 miles equal 30*1,6 = ? km step by step

jhonyy9:

@justus this mean 48 km / hour so than travel 48 km / hour in 2 hours how many km will travel ?

justus:

where did the 48 come from?

jhonyy9:

how ? than 48 km/h in 2 hours how you get halfed 48 and not doubled ?

jhonyy9:

30 miles per hour yes ? and given that 1 mile = 1,6 km so than 30 miles = ? km

justus:

oh I thought you meant to divide 48 by 2.

Imagine:

Justus use this.

jhonyy9:

@justus i think you made here a big mistake pls correct it

justus:

okay I had it right at first but the calculation went wrong

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus Did someone say cars? jk lol 1 mile is 1.6 km The rate is 30 km The rate (30 x 1) / 1.6 = \(\color{#0cbb34}{\text{End of Quote}}\)

Imagine:

\[30 miles = ? km\]

justus:

@artlover03 are you still there? :)

justus:

sorry the calculation was incorrect

jhonyy9:

no bc. you need multiplie the 30 by 1,6 not divide it

justus:

oh I see

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 no bc. you need multiplie the 30 by 1,6 not divide it \(\color{#0cbb34}{\text{End of Quote}}\) Correct.

justus:

so then yes it will be 48 * 2

jhonyy9:

MORE ATTENTION please

justus:

because we need to find the time it travels in 2 hours

jhonyy9:

indifferent when you ve converted 30 miles in km you made it incorrect

Imagine:

First you are trying to see how fast it goes per hour.

Imagine:

you skipped that part.

Imagine:

"What is the car's rate in kilometers per hour? "

justus:

The rate (30 x 1) * 1.6 = 48

justus:

48 *2 = 96

Imagine:

Correct. Yeah, that's right.

artlover03:

Im confused 😅

jhonyy9:

@justus pls explain it now

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 Im confused 😅 \(\color{#0cbb34}{\text{End of Quote}}\) sorry for the confusion it is 96

artlover03:

which is..?

Imagine:

\[(30 \times 1) * 1.6 = 48\] \[48 \times 2= 96km/hr\]

Imagine:

You are then trying to figure out how fast it goes in 2 hours.

Imagine:

"How many kilometers will the car travel in 2 hours?"

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine \[(30 \times 1) * 1.6 = 48\] \[48 \times 2= 96km/hr\] \(\color{#0cbb34}{\text{End of Quote}}\) 96 km/h not is correct

Imagine:

It has to be correct.

jhonyy9:

96 km in 2 hours

Imagine:

Yes.

justus:

I believe that is the final answer

artlover03:

so the rate is 48? and the distance traveled in 2 hours is 96

Imagine:

48 km/h is for one hour 96 km/h is for two hours.

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 so the rate is 48? and the distance traveled in 2 hours is 96 \(\color{#0cbb34}{\text{End of Quote}}\) yes

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus \(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 so the rate is 48? and the distance traveled in 2 hours is 96 \(\color{#0cbb34}{\text{End of Quote}}\) yes \(\color{#0cbb34}{\text{End of Quote}}\) I as well agree.

jhonyy9:

yes this is correct that 96 km travle in 2 hours but how @Imagine wrote there 96km/h so in this way wrote not is correct

Imagine:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine \[(30 \times 1) * 1.6 = 48km/h\] \[48 \times 2= 96km/h\] \(\color{#0cbb34}{\text{End of Quote}}\) It is correct. That's why the ending number is 48 in one hour, and 96 in 2.

justus:

okay now sorry for the inconvenience @artlover03. the error has been corrected. have a great day and good luck !!

jhonyy9:

@Imagine this 48*2 = 96 km without h

Imagine:

oh, correct.

artlover03:

Its all good! thank you for all your help! Its one of my final questions, i will post if its right or not soon :)

Imagine:

\[(30 \times 1) * 1.6 = 48km/h\] \[48 \times 2= 96km\]

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 Its all good! thank you for all your help! Its one of my final questions, i will post if its right or not soon :) \(\color{#0cbb34}{\text{End of Quote}}\) :) :)

jhonyy9:

bc. than you write 96 km/h this mean 96 km per hour

Imagine:

Correct, thanks!

justus:

Have a nice day everybody !

Imagine:

\[(30 \times 1) * 1.6 = 48km/h\] \[48 \times 2= 96km\] Final Answer.

jhonyy9:

correct now

Imagine:

Thanks Jhonny!

jhonyy9:

yw pls more attention

Imagine:

Will make sure of it.

jhonyy9:

and to @justus too

jhonyy9:

ok i will try being here to check it